OpenStudy (anonymous):

Why does ln((sinx)^1/2)=1/2ln(sinx)?

OpenStudy (anonymous):

\[\ln((sinx)^{1/2}) = 1/2\ln(sinx)\]

OpenStudy (anonymous):

what rule is being used here/when can it be used to pull the exponent out front?

OpenStudy (anonymous):

if it matters its a function i'm suposed to differentiate and my book says it would be easier if we simplified the function like that

OpenStudy (anonymous):

because it is the definition of logarithm that ln a^b = b ln a

OpenStudy (anonymous):

log 1000 = log 10^3 = 3 log 10...this is how log operates

OpenStudy (anonymous):

oh, ok thanks. Didn't know that

OpenStudy (raden_zaikaria):